40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 16:46:31 2019, gautam, Update, IOO, IMC not locking PCdrive_RMS.png
    Reply  Tue Jun 4 10:56:31 2019, gautam, Update, IOO, IMC diagnostics IMCdiag.pdf
       Reply  Thu Jun 6 22:11:53 2019, Koji, Update, IOO, IMC diagnostics 
          Reply  Thu Jun 13 18:01:18 2019, aaron, Update, IOO, IMC diagnostics 
             Reply  Fri Jun 14 13:10:00 2019, aaron, Update, IOO, IMC diagnostics 
                Reply  Mon Jun 17 12:37:16 2019, aaron, Update, IOO, IMC diagnostics 
                   Reply  Tue Jun 18 20:35:07 2019, aaron, Update, IOO, IMC diagnostics 
                      Reply  Wed Jun 19 19:12:51 2019, aaron, Update, IOO, IMC diagnostics PMC_Error_Spectrum.pdfPDH_signal.pdfPDH_signal_full.pdf
                         Reply  Mon Jun 24 15:49:05 2019, aaron, Update, IOO, IMC diagnostics sweep_config_updated.png
                            Reply  Tue Jul 2 19:36:18 2019, aaron, Update, IOO, IMC diagnostics PDH_error.pdfPMC_Error_Spectrum.pdf
          Reply  Fri Jun 21 19:36:26 2019, Koji, Update, IOO, IMC diagnostics IMC_error.pdf
             Reply  Sun Jun 23 08:09:53 2019, gautam, Update, IOO, NPRO diagnostics consolidatedOscopeScreenCaps.pdfRINcomp.pdfPLL_OLTF.pdfPLLnoise.pdf
                Reply  Sun Jun 23 09:36:32 2019, gautam, Update, IOO, IMC is locking normally again PMCdiag.pdf
                   Reply  Sun Jun 23 14:43:14 2019, Koji, Update, IOO, IMC is locking normally again 
                   Reply  Mon Jun 24 08:12:10 2019, gautam, Update, IOO, IMC is locking normally again IMCdutycycle.png
                      Reply  Mon Jun 24 11:48:35 2019, gautam, Update, IOO, IMC in-loop error spectra and OLTF IMCfreqNoise.pdfIMC_OLTF.pdfIMC_CMboard.jpg
Message ID: 14659     Entry time: Thu Jun 6 22:11:53 2019     In reply to: 14653     Reply to this: 14670   14686
Author: Koji 
Type: Update 
Category: IOO 
Subject: IMC diagnostics 

As per Gautam's request, I looked at the IMC situation.

Locking path

  • Acquisition: IMC IN1 Gain +4 (nominal), Boost 0, VCO Gain (-32), FSS Common +6 (nominal), FSS FAST +20
    This is too low gain. So oscillate VCO Gain between -32 and ~0 until TEM00 lock is acquired
  • Once lock is acquired, bring the VCO gain to +11 (new nominal), and increase the FSS FAST to +23 (new nominal). Change the IMC BOOST to 3 (nominal)

Diagnosis

  • The PMC servo gain was checked. The control signal monitor for the PMC actuation was hooked up to SR785. The nominal gain was +18dB. Increasing the gain to 20dB made the servo oscillating. So the nominal gain of +18dB seems still reasonable.
  • The status of NOISE EATER was checked. Both the PMC REFL and TRANS were looked at by AG4395A. The power spectrum of them did not change much around the kHz~MHz region. It made the PSD slightly (x2~3) improved below 1kHz. I also did not recognize the relaxation oscillation peak. So I could not figure out where to see. NOISE EATSER was on and is still on.
  • IFO Modulation Freq: I took this chance to look at the IMC absolute length using the peak at 3.6MHz. The TP1A output of the IMC servo board was hooked up to AG4395A.
    The new FSR of the IMC (and thus the modulation frequency for the IFO) is 11.066275MHz (instead of the previous 11.066209MHz).
    This corresponds to 0.16mm difference in the roundtrip length.
  • (*Still working) IMC SERVO configuration:
    • FAST 25 (nominal) sometimes invoke the oscilattion. 24 has gain peaking ~30kHz. There is a big line peak at 35kHz so wanted to avoid the servo bump (PZT-EOM cross over). So decided to use 23dB. (This is not optimal for the CM servo as we need as much as bandwidth for CM servo.)
    • IMC VCO GAIN (bad name. this is actually overall output gain for IMC) was increased from the nominal 7 to 11. Increasing this above 11 makes the servo oscillating at ~200kHz.
  • (*Still working) Measured power spectrum of the error signal. Too many line peaks.
  • (*Still working) Single trigger observation: Oscilloscope monitoring started from 35kHz going up and ~20kHz oscillation +/-6V of the IMC servo output was observed. Could not capture good data for this. Try the other day.

I'll complete the entry later.

ELOG V3.1.3-